OpenStudy (anonymous):

What is the equation for the line that passes through the coordinates (1,-2), (2,-6) 1.) y=4x+2 2.) y+4x=2 3.) 4y-x=2 4.) 4y+x=2

OpenStudy (zzr0ck3r):

want to know why?

OpenStudy (anonymous):

x + 4y=2

OpenStudy (anonymous):

Yes i do! I have another question just like it, and I really cant figure it out!

OpenStudy (zzr0ck3r):

what is the slope of these two points?

OpenStudy (zzr0ck3r):

\[\frac{y_1-y_0}{x_1-x_0}\]

OpenStudy (anonymous):

3/2?:o

OpenStudy (zzr0ck3r):

(1,-2), (2,-6) so \[\frac{-6-(-2)}{2-1}=?\]

OpenStudy (anonymous):

\[\frac{ -4 }{ 1 }\]

OpenStudy (zzr0ck3r):

=-4

OpenStudy (zzr0ck3r):

right

OpenStudy (anonymous):

yeah

OpenStudy (zzr0ck3r):

ok so do you get how I did that with \[\frac{y_1-y_0}{x_1-x_0}\] ?

OpenStudy (zzr0ck3r):

makes sense so far?

OpenStudy (zzr0ck3r):

we found the slope

OpenStudy (anonymous):

yeah it does

OpenStudy (zzr0ck3r):

ok so do you know this y=mx+b?

OpenStudy (zzr0ck3r):

have you seen this before?

OpenStudy (anonymous):

Isnt that point slope form?

OpenStudy (zzr0ck3r):

slope intercept

OpenStudy (anonymous):

oh okay! By the way if i am coming off as ignorant I'm sorry, math has just never really stuck with me too well

OpenStudy (zzr0ck3r):

ok so we know the value for m right? m=-4 so y=-4x+b and all we need is to find out what b is. Make sense?

OpenStudy (anonymous):

Yeah

OpenStudy (zzr0ck3r):

np we all know only what we have learned:)

OpenStudy (zzr0ck3r):

ok so y=-4x+b we can use either point to find out what b is (because both points will lie on our line) we will use (2,-6) so in this point x=2 and y=-6 so plug this into what we have (y=-4x+b) so -6=-4(2)+b make sense?

OpenStudy (anonymous):

Yeah I think i got it

OpenStudy (anonymous):

does 2=b or am I way ahead of myself?

OpenStudy (anonymous):

y=4x+2 would be my final answer ?

OpenStudy (zzr0ck3r):

well b=2 correct but remember slope was not 4 it was -4 so y=-4x+2

OpenStudy (anonymous):

Now I'm confused:o

OpenStudy (anonymous):

Because that isnt one of my possible answers:o

OpenStudy (zzr0ck3r):

lol sorry I was helping someone else.

OpenStudy (anonymous):

Ahah it's fine!

OpenStudy (zzr0ck3r):

ahh but it is, it just does not look like one y=-4x+2 add 4x to both sides and tell me what you have.

OpenStudy (anonymous):

would it be 4y+x=2??

OpenStudy (zzr0ck3r):

y=-4x+2 add 4x to both sides of this

OpenStudy (zzr0ck3r):

they gave you the wrong answer...

OpenStudy (anonymous):

Really??

OpenStudy (zzr0ck3r):

yes really x + 4y=2 take one of your points (1,-2) x=1 and y=-2 plug into the one they said x + 4y=2 1+4(-2)=2 -7=2

OpenStudy (anonymous):

Because, the program I am using is a summer school program to make up a credit for a class I failed by half a percentage point xb And it's terrible for teaching me any of the subjects needed

OpenStudy (zzr0ck3r):

ok back to this problem y=-4x+2 add 4x to both sides of this y+4x=-4x+2+4x=-4x+4x+2=2 so y+4x=2

OpenStudy (anonymous):

Ahah, so then if that is the final answer, i should stick with D as my answer on the program?

OpenStudy (zzr0ck3r):

ok this is what we did we were given two points (x_1,y_1) and (x_0,y_0) we need y=mx+b slope = \[\frac{y_1-y_0}{x_1-x_0}=slope=m\] so we know m y=mx+b use either point and plug in the x value for x and the y value for y then solve for b now we know m and b y=mx+b

OpenStudy (zzr0ck3r):

yes d

OpenStudy (zzr0ck3r):

wait

OpenStudy (anonymous):

Okay I'm pretty sure i get it now, or atleast a hell of a lot more than I ever did before. I'll try the same steps on the last question that I have just like this

OpenStudy (zzr0ck3r):

no y+4x=2

OpenStudy (zzr0ck3r):

this is option2)

OpenStudy (zzr0ck3r):

they told you option 4, and that was wrong we proved option 2 was right

OpenStudy (anonymous):

Ahaha, alright thank you!:D Hopefully I can do the last problem and be done with this test now!